# Adaptive Hyper-Graph Convolution Network for Skeleton-based Human Action Recognition with Virtual Connections

### v1

@@ Abstract
The shared topology of human skeletons motivated the recent investigation of graph convolutional network (GCN) solutions for action recognition. However, the existing GCNs rely on the binary connection of two neighbouring vertices (joints) formed by an edge (bone), overlooking the potential of constructing multi-vertex convolution structures. In this paper we address this oversight and explore the merits of a hyper-graph convolutional network (Hyper-GCN) to achieve the aggregation of rich semantic information conveyed by skeleton vertices. In particular, our Hyper-GCN adaptively optimises multi-scale hyper-graphs during training, revealing the action-driven multi-vertex relations. Besides, virtual connections are often designed to support efficient feature aggregation, implicitly extending the spectrum of dependencies within the skeleton. By injecting virtual connections into hyper-graphs, the semantic clues of diverse action categories can be highlighted. The results of experiments conducted on the NTU-60, NTU-120, and NW-UCLA datasets, demonstrate the merits of our Hyper-GCN, compared to the state-of-the-art methods. Specifically, we outperform the existing solutions on NTU-120, achieving 90.2% and 91.4% in terms of the top-1 recognition accuracy on X-Sub and X-Set.
@@ I Introduction
Skeleton-based human action recognition is a popular research topic in artificial intelligence, with practical applications in video understanding, video surveillance, human-computer interaction, robot vision, VR, and AR [1, 2, 3, 4, 5, 6, 7]. In general, a skeleton sequence contains a series of 2D or 3D coordinates, which can easily be collected by low-cost depth sensors or obtained by video-based pose estimation algorithms. Compared to RGB and optical flow images, skeleton data, which represents the basic physical structure of a human being, is of lower dimension and higher efficiency. Moreover, it is robust to illumination changes and scene variations. For these reasons, the adoption of this structural data is very popular in skeleton-based action recognition [8, 9, 10, 11, 12, 13, 14].
@@
Figure 1: Illustration of our Hyper-GCN. In (a), dotted lines represent the virtual connections. Each coloured part containing multiple joints represents a hyper-graph with hyper-edges. (b) provides an intuitive comparison between normal graph convolution and hyper-graph convolution operations with the same connection degree.
@@
To facilitate skeleton-based action recognition, both Recurrent Neural Networks (RNNs) [8, 9, 10, 11] and Convolutional Neural Networks (CNNs) [15, 16, 17] have been well explored. However, RNNs themselves cannot depict the intrinsic skeleton topology. The learned filters of CNNs, on the other hand, neglect the spatio-temporal structure of the skeleton. Drawing on these observations, recent studies have focused on how to model the skeleton topology directly.
@@
In principle, the physical topology of human joints and bones can be consistently represented by a graph. Accordingly, the graph convolutional network (GCN) is typically introduced to achieve the aggregation of feature information conveyed by skeleton joints [18, 13, 19]. In general, two neighbouring joints can pass messages through their shared bone. In graph terms, this corresponds to the exchange of information between two vertices along their connecting edge. Besides the physical skeleton topology, attempts have been made to explore implicit relationships among joints [20, 13, 12, 19], suggesting different variants of skeleton topology. In principle, these approaches assume a binary connection between each connected vertex pair. Mathematically, the constructed topology in the adjacency matrix is represented by a normal graph. However, human actions are jointly defined by several joints. Hence, human actions encompass not only binary relations between vertex pairs but also multi-joint relationships. For instance, the action primitive starting running is manifest in the human raising the left hand while the right leg steps forward, as shown in Figure 1 (a). The binary connections are not sufficient to capture the synergistic interaction of multiple joints. This strongly argues for constructing feature aggregation paths involving multiple vertices.
@@
Accordingly, we propose to construct a hyper-graph to depict the skeleton topology and take advantage of the outstanding performance of hyper-graph analysis techniques  [21, 22, 23, 24]. The hyper-graph topology involves multiple node connections, rather than binary connections of a normal graph. As shown in Figure 1 (a), in a hyper-graph, a hyper-edge can link more than two vertices, which has the capacity to represent complex collaborative relations among human joints. As one hyper-edge associates multiple vertices, a single hyper-graph convolution enables aggregating all the features along the hyper-edges. An illustration is provided in Figure 1 (b), where the normal graph and hyper-graph are presented to demonstrate their differences in passing the information conveyed by the vertex features.
@@
Conceptually, for a fair comparison, we typically set each hyper-edge to contain only 2 vertices, which is the same as the normal graph. In the normal graph convolution, after 2-layer aggregation, the information of the green vertex is spread to 2 other vertices. In contrast, in a hyper-graph convolution, the information of the green vertex spreads to 4 vertices including itself, creating an extended receptive field. Theoretically, by applying a hyper-graph for modelling joint aggregation on a skeleton, improved efficiency can be obtained during learning the action semantics.
@@
Besides its information aggregation mode, the capacity of an action recognition system is also limited by its input features. In general, the number of skeleton joints is fixed in existing benchmarks, e.g., 25 for NTU-120 [25]. The entire recognition process relies on the interactions among the involved joints. In the domain of artistic puppetry, the history of driving actions, such as Shadow Play  [^1] https://en.wikipedia.org/wiki/Shadow_play and Marionette  [^2] https://en.wikipedia.org/wiki/Marionette goes back more than 2000 years. This kind of art form provides an inspiration for involving additional ’hyper joints’ which can drive or facilitate a better communication between existing joints. The underlying spirit is to alleviate the pressure of the real joints in storing and transferring semantics. Jointly with the hyper joints, real joints can focus more on storing neighbouring joint features and transferring the global information to the hyper joints. Interestingly, the class token in existing Transformers can also be considered as a hyper-token  [26, 27, 28, 29]. As shown in Figure 1, the skeleton is described as a marionette, where the actions are ”controlled” by connecting hyper joints to the real joints. This suggests that the hyper joints are not only able to capture the representation information of human action, but also reveal the implicit information between physically connected joints as hints for recognition.
@@
By endowing a hyper-graph with hyper joints, virtual connections are created to perform comprehensive hyper-graph convolutions. We construct Hyper-GCN based on the above design principles. Extensive experiments on 3 datasets, NTU RGB+D 60, NTU RGB+D 120, and NW-UCLA, are conducted for evaluation. The results validate the merits of our proposed Hyper-GCN. The main contributions are as follows:
  • We propose to represent human skeleton topology by a hyper-graph. Compared with the normal graph, the efficiency of feature interaction is improved via multi-vertex aggregation.
  • Virtual hyper joints are injected to connect the physical joints, enhancing the model capacity to capture the global semantics.
  • We propose an adaptive hyper-graph convolution network (Hyper-GCN) for skeleton-based human action recognition. On all three public datasets used for experimentation, Hyper-GCN achieves the SOTA performance.
@@ II-A Graph Topology for Action Recognition
A graph can represent a human skeleton, preserving the joint relationships via predefined edges. To aggregate semantics, GCNs have been well studied in recent years [30, 31, 32, 33, 34, 35, 36, 37, 38]. Regarding modelling the spatio-temporal relevance, ST-GCN [18] proposes to represent the topology with an adjacency matrix of 3 subsets. Similarly, 2s-AGCN [20], InfoGCN [12], and DS-GCN [14] use a self-attentive mechanism to learn the topology from joint pairs. Besides the intuitive spatio-temporal dimensions, CTR-GCN [13] proposes to learn the channel topology using subspace mappings. All the above methods construct the topology from joint pairs, which rely only on the binary relation between two vertices. In their implementation stage, the adjacency matrix is used to represent the topology. In this case, high-order information among joints is not taken into consideration, neglecting the collaborative power among multiple joints. Though hyper-graph is considered to construct topology using multivariate joint relationships, current [39, 40] solutions manually set the hyper-graphs, which greatly relies on human experience, sacrificing the adaptability of graph learning.
II-B Feature Settings for Action Recognition
It has been observed that the input features play the most essential role in delivering high recognition accuracy [41, 42, 43, 36, 44]. Drawing on this, Graph2Net [45] proposes to extract local and global spatial features jointly. Similarly, CTR-GCN [13] uses multi-scale temporal convolution to extract temporal features. While HD-GCN [19] focuses on learning the feature aggregation among different edge sets. To extend the perception field, STC-Net [46] uses the dilated kernels for graph convolution to capture the features. After all, existing GCNs [18, 20, 12, 14, 47] receive the input features from real skeleton joints. However, each skeleton joint acts as an information carrier during forward passing, which is required to deliver both local context and global semantics. Given a fixed representative capacity (number of channels), we believe it is necessary to involve additional virtual joints to balance the pressure of storing local and global information.
@@ III-A Preliminaries
$\mathbf{F_{out}}=\sigma(\hat{A}\mathbf{F_{in}}W),$ (2)
where $W\in\mathbb{R}^{C\times C^{\prime}}$ is the learnable parameters, representing the feature transformation patterns in the feature space. $\sigma$ denotes the non-linear activation function ReLU.
@@ III-B Hyper-graph
III-C Adaptive Hyper-graph Construction Module
How to construct the hyper-graph is critical to better understanding human actions. As shown in Figure 2 (b), we design an Adaptive Hyper-graph Construction Module (AHC-Module). Given the feature $\mathbf{{F}^{{}^{\prime}}_{in}}\in\mathbb{R}^{C\times T\times V}$, we first perform temporal pooling, and then the obtained spatial information is the evidence to determine the optimal hyper-graph construction. This can efficiently decouple the temporal and spatial clues. After performing temporal average-pooling, we can obtain $\mathbf{\bar{F}^{{}^{\prime}}_{in}}\in\mathbb{R}^{C\times V}$. To adjust the scale of the input feature channels, we adopt a layer normalisation operation on the channel dimension. After that, we use two individual mapping functions to embed the features into the subspace for constructing the hyper-graph. This design enables preserving the original features to reflect their spatial relevance. The embedding operation can be formulated as follows:
$P_{\textrm{emb}}$ $=\Phi(\textrm{LN}(\textrm{Pooling}_{\textrm{T}}(\mathbf{F^{{}^{\prime}}_{in}}))),$ (7)
$W_{\textrm{emb}}$ $=\textrm{Tanh}(\Psi(\textrm{LN}(\textrm{Pooling}_{\textrm{T}}(\mathbf{F^{{}^{\prime}}_{in}})))),$ (8)
where $P_{\textrm{emb}}$ represents the relative position of the joints in the embedding space, and $W_{\textrm{emb}}$ represents the weights of hyper-edges. $\textrm{Pooling}_{\textrm{T}}$ denotes performing pooling in the temporal dimension. LN denotes the layer normalisation along the channel dimension. $\Phi\in\mathbb{R}^{C\times C_{h}},\Psi\in\mathbb{R}^{C\times 1}$ are the mapping functions of position and weight, which are set learnable parameters. $C_{h}$ represents the channels of the mapping subspace. Tanh denotes the activation function to obtain the weight of each hyper-edge, limiting the values with the range of $[-1,1]$.
Figure 3: Illustration of the hyper-graph construction process with the proposed AHC-Module. $N$ represents the number of joints.
In the embedding space, our AHC-Module utilises the Euclidean distance to measure the joint difference. For the mapping of hyper-edges and vertices, the reason we apply only one mapping function instead of two separate ones is to ensure that vertices and hyper-edges are within the same semantic space. Define the set of joints $P_{\textrm{emb}}=\{v_{0}\in\mathbb{R}^{C_{h}},v_{1}\in\mathbb{R}^{C_{h}},...,v_{N}\in\mathbb{R}^{C_{h}}\}$, given $v_{i},v_{j}\in P_{\textrm{emb}}$, each element $m_{i,j}$ in the distance matrix $M\in\mathbb{R}^{N\times N}$ can be obtained as follows:
$dis_{i,j}=dis_{j,i}=||v_{i}-v_{j}||^{2}_{2}.$ (9)
@@ III-C Adaptive Hyper-graph Construction Module
In principle, we need to transform the distance matrix into a probability hyper-grap0h, which determines whether the specific joints are contained with the same hyper-edges. In order to guarantee that the model is trainable, we use the softmax operation to assign the probabilities. However, directly using the softmax would inevitably result in each joint belonging to all hyper-edges to some extent. Since this outcome is undesirable, we introduce constraints on each joint to limit the number of hyper-edges it can belong to, preventing excessive connections. We retain only the $K$ nearest hyper-edges represented by each joint in the semantic space for probability estimation. Given the row vector $\mathbf{m}_{i}$ in the distance matrix $M$, which represents the distances between joint $i$ and the other joints, we select the indices of the minimum $K$ joints from $\mathbf{m}_{i}$ and add them to the set $set^{K}_{i}$. Based on $set^{K}_{i}$, $h_{i,j}$ in the incidence matrix $H\in\mathbb{R}^{N\times N}$ of hyper-graph, can be calculated as follows:
$h_{i,j}$ =$=\begin{cases}\frac{e^{-m_{i,j}}}{\sum_{k\in set^{K}_{i}}{e^{-m_{i,k}}}},&j\in set^{K}_{i},\\ 0,&j\notin set^{K}_{i}\end{cases}$ (10)
The shape of our incidence matrix is consistent with the adjacency matrix of the normal graph. While our topology with hyper-graph contains more high-order information. The entire process is illustrated in the Figure 3.
III-D Multi Scale Hyper-graph Convolution
As our AHC-Module can depict the topology into a hyper-graph at different scales by adjusting the value of $K$. To accommodate the semantic information reflected by these varying scales, we further propose the Multi-Scale Hyper-graph Convolution (MS-HGC), as shown in Figure 2 (a). We use different branches to independently perform hyper-graph convolution on the topological information represented by the multi-scale hyper-graphs, thereby enhancing the computational efficiency of the Hyper-GCN.
Besides the hyper-graph, we incorporate the physical topology to emphasise the natural physical relations of a human being’s skeleton. To achieve this objective, existing methods [18, 20, 13, 19] divide the physical topology of the human body into 3 subsets. They are represented by $S=\{s_{id},s_{cf},s_{cp}\}$, where $s_{id}$, $s_{cf}$ and $s_{cp}$ denote the identity, centrifugal, and centripetal joint subsets. In our MS-HGC, to ensure that the integrated physical topology in each branch remains complete, we merge them into a single set. We divide the input feature into 8 separate branches along the channel dimension, which are processed by 8 MS-HGC parallelly, delivering 8 multi-scale hyper-graphs. 8 branches are aggregated by channel concatenation after the hyper-graph convolutions, as shown in Figure 2 (a). The operation of MS-HGC can be formulated as follows:
$\mathbf{F_{out}}$ $=\uplus^{8}_{i=1}{(\hat{A_{i}}+\alpha_{i}\cdot\hat{H_{i}})\mathbf{F^{[(i-1)\times C^{{}^{\prime}}+1,...,i\times C^{{}^{\prime}}]}_{in}}W_{i}},$ (11)

### v2

@@ Abstract
The shared topology of human skeletons motivated the recent investigation of graph convolutional network (GCN) solutions for action recognition. However, most of the existing GCNs rely on the binary connection of two neighboring vertices (joints) formed by an edge (bone), overlooking the potential of constructing multi-vertex convolution structures. Although some studies have attempted to utilize hyper-graphs to represent the topology, they rely on a fixed construction strategy, which limits their adaptivity in uncovering the intricate latent relationships within the action. In this paper, we address this oversight and explore the merits of an adaptive hyper-graph convolutional network (Hyper-GCN) to achieve the aggregation of rich semantic information conveyed by skeleton vertices. In particular, our Hyper-GCN adaptively optimises the hyper-graphs during training, revealing the action-driven multi-vertex relations. Besides, virtual connections are often designed to support efficient feature aggregation, implicitly extending the spectrum of dependencies within the skeleton. By injecting virtual connections into hyper-graphs, the semantic clues of diverse action categories can be highlighted. The results of experiments conducted on the NTU-60, NTU-120, and NW-UCLA datasets demonstrate the merits of our Hyper-GCN, compared to the state-of-the-art methods. Specifically, we outperform the existing solutions on NTU-120, achieving 90.5% and 91.7% in terms of the top-1 recognition accuracy on X-Sub and X-Set.
@@ I Introduction
Skeleton-based human action recognition is a popular research topic in artificial intelligence, with practical applications in video understanding, video surveillance, human-computer interaction, robot vision, VR, and AR [1, 2, 3, 4, 5, 6, 7]. In general, a skeleton sequence contains a series of 2D or 3D coordinates, which can easily be collected by low-cost depth sensors or obtained by video-based pose estimation algorithms. Compared to RGB and optical flow images, skeleton data, which represents the basic physical structure of a human being, is of lower dimension, conveying human action with higher efficiency. Moreover, it is robust to illumination changes and scene variations. For these reasons, the adoption of this structural data is very popular in skeleton-based action recognition [8, 9, 10, 11, 12, 13, 14].
@@
Figure 1: Illustration of our Hyper-GCN. In (a), dotted lines represent the virtual connections. Each coloured part containing multiple joints represents a hyper-graph with hyper-edges. (b) provides an intuitive comparison between normal graph convolution and hyper-graph convolution operations with the same connectivity degree.
@@
To facilitate skeleton-based action recognition, both Recurrent Neural Networks (RNNs) [8, 9, 10, 11] and Convolutional Neural Networks (CNNs) [15, 16, 17] have been well explored. However, RNNs themselves cannot depict the intrinsic skeleton topology. The learned filters of CNNs, on the other hand, neglect the spatio-temporal structure of the skeleton. Drawing on these observations, recent studies have focused on how to model the skeleton topology directly. In principle, the physical topology of human joints and bones can be consistently represented by a graph. Accordingly, the graph convolutional network (GCN) is typically introduced to aggregate feature information conveyed by skeleton joints [18, 19, 20]. In terms of data representations, skeleton joints are also one-dimensional sequence data, similar to language. Therefore, with the rise of recent self-attention techniques [21, 22], several Transformer-based action recognition solutions [23, 24, 25, 26] have been released. Transformer considers skeletal joints as tokens and uses attention maps to reflect the topology. These methods have achieved excellent performance but sacrifice much higher GFLOPS and parameters than GCN.
@@
In general, two neighbouring joints can pass messages through their shared bone. In graph terms, this corresponds to the exchange of information between two vertices along their connecting edge. Besides the physical skeleton topology, attempts have been made to explore the implicit relationships among joints [27, 13, 12, 19], suggesting different variants of skeleton topology. In principle, these approaches assume a binary connection between each connected vertex pair. Mathematically, the constructed topology in the adjacency matrix is represented by a normal graph. However, human actions are jointly defined by several joints. Hence, human actions encompass not only binary relations between vertex pairs but also multi-joint relationships. For instance, the action primitive starting running is manifest in the human raising the left hand while the right leg steps forward, as shown in Figure 1 (a). The binary connections are not sufficient to capture the synergistic interaction of multiple joints. This strongly argues for constructing feature aggregation paths involving multiple vertices.
@@
Accordingly, we propose to construct a hyper-graph to depict the skeleton topology and take advantage of the outstanding performance of hyper-graph analysis techniques  [28, 29, 30, 31]. The hyper-graph topology involves multiple node connections, rather than binary connections of a normal graph. As shown in Figure 1 (a), in a hyper-graph, a hyper-edge can link more than two vertices. This linking has the capacity to represent complex collaborative relations among human joints. As one hyper-edge associates multiple vertices, a single hyper-graph convolution enables aggregating all the features along the hyper-edges.
@@
An illustration is provided in Figure 1 (b), where the normal graph and hyper-graph are presented to demonstrate their differences in passing the information conveyed by the vertex features. In the normal graph convolution, after 2-layer aggregation, the information of the green vertex is spread to 2 other vertices. In contrast, in a hyper-graph convolution, the information of the green vertex spreads to all the vertices, creating an extended receptive field. Theoretically, by modelling a skeleton using a hyper-graph for the aggregation of joint information, improved efficiency can be obtained during learning the action semantics.
However, the key to achieving enhanced information aggregation by hyper-graph convolution lies in the conformity of the hyper-graph structure to the real topology of human action. Several studies [32, 33] attempt to adopt a hyper-graph to model the skeleton topology, mostly based on superficial prior knowledge and an unfounded definition of the structure of hyper-graph. Distinguishing our approach from the existing hyper-graph-based methods, we design an adaptive solution for constructing the hyper-graph with virtual connections. It can construct a Non-uniform hyper-graph specific to each individual action type.
@@
Besides the information aggregation issue, the capacity of an action recognition system is also constrained by its input features. In general, the number of skeleton joints is fixed in existing benchmarks, e.g., 25 for NTU-120 [34]. The entire recognition process relies on a successful capture of the interactions among the skeleton joints. In the domain of artistic puppetry, the history of driving actions, such as Shadow Play  [^1] https://en.wikipedia.org/wiki/Shadow_play and Marionette  [^2] https://en.wikipedia.org/wiki/Marionette goes back more than 2000 years. This kind of art form provides an inspiration for involving additional ’hyper joints’ which can drive or facilitate a better communication between existing joints. The underlying spirit is to alleviate the pressure on the real joints to store and transfer complex semantics. Jointly with the hyper joints, real joints can focus more on storing neighbouring joint features and relegating the task of transferring the global information to the hyper joints. Interestingly, the class token in existing Transformers can also be considered as a hyper-token  [21, 22]. As shown in Figure 1, the skeleton is described as a marionette, where the actions are ”controlled” by connecting hyper joints to the real joints. This suggests that the hyper joints are not only able to capture the representation information of human action, but also reveal the implicit information between physically connected joints as hints for recognition.
@@
By endowing an adaptive non-uniform hyper-graph with hyper joints, virtual connections are created to perform comprehensive hyper-graph convolutions. We construct Hyper-GCN based on the above design principles. Extensive experiments on 3 datasets, NTU RGB+D 60 [11], NTU RGB+D 120 [34], and NW-UCLA [35], are conducted for evaluation. The results validate the merits of our proposed Hyper-GCN. The main contributions are as follows:
  • An adaptive non-uniform hyper-graph to represent the human skeleton topology. Compared with a fixed hyper-graph, the constructed topology is more action-specific, thus boosting the discrimination.
  • The injection of virtual hyper joints, enriching the connectivity of the physical joints, from a global semantic perspective.
  • As the processing architecture, we propose a hyper-graph convolution network (Hyper-GCN). The SOTA performance achieved on three public datasets demonstrates the merits of the Hyper-GCN and virtual connection designs.
@@ II-A Graph Topology for Action Recognition
A graph can represent a human skeleton, preserving the joint relationships via predefined edges. To aggregate semantics, GCNs [36, 37, 38, 39, 40, 41, 42, 20] and Transformers  [23, 43, 25, 24, 44] have been well studied.
For GCN-based approaches, ST-GCN [18] proposes to represent the topology with an adjacency matrix of 3 subsets by modeling the spatio-temporal relevance. Similarly, 2s-AGCN [27], InfoGCN [12], and DS-GCN [14] use a self-attentive mechanism to learn the topology from joint pairs. Besides the intuitive spatio-temporal dimensions, CTR-GCN [13] proposes to learn the topology for channels to refine the skeleton.
For Transformer-based solutions, IIP-Transformer [26] adopt a self-attention mechanism to establish the intra-part and the inter-part joints. IGFormer [45] learns the topology between persons at both semantic and distance levels. STTFormer [46] proposed structure can capture the correlation between joints in consecutive frames.
All the above two paradigms construct the topology from joint pairs, which rely only on the binary relation between two vertices. In their implementation stage, the adjacency matrix or attention map is used to represent the topology. In this case, high-order information among joints is not taken into consideration, neglecting the collaborative power among multiple joints. Though hyper-graph is considered to construct topology using multivariate joint relationships, current [33, 32] solutions manually set the hyper-graphs, which greatly relies on human experience, sacrificing the adaptability of graph learning.
II-B Feature Configurations for Action Recognition
It has been observed that the input features play the most essential role in delivering high recognition accuracy [47, 48, 49, 42, 50, 26, 45, 46] in GCNs and Transformers.
Drawing on this, Graph2Net [51] proposes to extract local and global spatial features jointly. CTR-GCN [13] uses multi-scale temporal convolution to extract temporal features. While HD-GCN [19] introduce the hierarchical edge convolution to key edge features. To extend the perception field, STC-Net [52] uses the dilated kernels for graph convolution to capture the features.
Similarly, IG-Transformer [45] aggregate the features of two persons as one to mine the semantic information. STST [43] split the coordinate features, semantic features and temporal features into three kinds of tokens. TSTE [24] extracting and merging the spatial and temporal features as input to transformer encoders.
After all, existing methods [18, 27, 12, 14, 53, 25, 23] receive the input features from real skeleton joints. However, each skeleton joint acts as an information carrier during forward passing, which is required to deliver both local context and global semantics. Given a fixed representative capacity (number of channels), we believe it is necessary to involve additional virtual joints to balance the pressure of storing local and global information.
@@ III-A Preliminaries
$\mathbf{F_{out}}=\sigma(\hat{A}\mathbf{F_{in}}P),$ (2)
where $P\in\mathbb{R}^{C\times C^{\prime}}$ is the learnable parameters, representing the feature transformation patterns in the feature space. $\sigma$ denotes the non-linear activation function ReLU.
@@ III-B Hyper-graph
Figure 3: Instance of the incidence matrix construction by A-NHG with hyper-parameter $K=3$. $N$ represents the number of joints. (a) represents the hyper-edge sets are null. (b) represents that measure the distance between 1-st joint and other joints. (c) represents that retrain $K$-nearest hyper-edges which contain 1-st joint. (d) represents the final state.
III-C Adaptive Non-uniform Hyper-graph
Constructing the hyper-graph incidence matrix is critical to understanding human actions. As shown in Figure 2 (b), we design an Adaptive Non-uniform Hyper-graph (A-NHG) Construction.
Given the general feature $\mathbf{X}\in\mathbb{R}^{N\times C}$ of spatial skeleton joints, $N$ represent the number of joints, $C$ represent the channels. Our A-NHG construct the hyper-edge with each joint as the center of mass, and a total of $N$ hyper-edges are constructed. A-NHG utilises the Euclidean distance to measure the joint difference.
We use one mapping function $\Phi\in\mathbb{R}^{C\times C_{h}}$ to embed the features into the subspace for constructing the hyper-graph. $C_{h}$ represents the hidden channels of the mapping subspace. This design enables preserving the original features to reflect their spatial relevance. Based on $X_{H}\in\mathbb{R}^{N\times C_{h}}$, we define the set of joints $\mathcal{V_{H}}=\{v_{1}\in\mathbb{R}^{C_{h}},v_{2}\in\mathbb{R}^{C_{h}},...,v_{N}\in\mathbb{R}^{C_{h}}\}$. Given $v_{i},v_{j}\in\mathcal{V_{H}}$, each element $m_{i,j}$ in the distance matrix $M\in\mathbb{R}^{N\times N}$ can be obtained as follows:
$m_{i,j}=m_{j,i}=||v_{i}-v_{j}||^{2}_{2}.$ (7)
@@ III-C Adaptive Non-uniform Hyper-graph
In principle, we need to transform the distance matrix into a probability incidence matrix, which determines whether the specific joints are contained with the same hyper-edges. In order to guarantee that the model is trainable, we use the softmax operation to assign the probabilities. However, directly using the softmax would inevitably result in each joint belonging to all hyper-edges to some extent.
Since this outcome is undesirable, we introduce constraints on each joint to limit the number of hyper-edges it can belong to, preventing excessive connections. For one joint, we retain only the $K$ nearest hyper-edges represented by joints in the semantic space for probability estimation. And $K$ is the hyper-parameter of A-NHG.
Given the row vector $\mathbf{m}_{i}$ in the distance matrix $M$, which represents the distances between joint $i$ and the other joints. We select the indices of the minimum $K$ joints from $\mathbf{m}_{i}$ and add them to the set $set_{i}$. Based on $set_{i}$, $h_{i,j}$ in the incidence matrix $H\in\mathbb{R}^{N\times N}$, can be calculated as follows:
$h_{i,j}$ =$=\begin{cases}\frac{\text{exp}(-m_{i,j})}{\sum_{k\in set_{i}}{\text{exp}(-m_{i,k})}},&j\in set_{i},\\ 0,&j\notin set_{i}\end{cases}$ (8)
In contrast to uniform hyper-graph (every hyper-edge contains the same number of joints), our approach limits the maximum number of every joint that can be contained by hyper-edges. Obviously, the number of joints contained in each hyper-edge constructed in this way is non-fixed, allowing the hyper-edges to capture more differentiated associations. The entire process for example with $K$ = 3 is illustrated in the Figure 3.
III-D Multi-head Hyper-graph Convolution
To accommodate the semantic information reflected by different group channels, we further propose the Multi-head Hyper-graph Convolution (M-HGC), as shown in Figure 2 (a).
We use separate branches to independently perform hyper-graph convolution on the topologies represented by the multi-head hyper-graphs, thereby enhancing the computational efficiency of the Hyper-GCN. For instance, the feature $\mathbf{F_{in}\in\mathbb{R}^{C_{in}\times T\times V}}$ is divided into $8$ separate branches along the channel dimension, which are processed by $8$ M-HGC parallelly, delivering $8$ separate hyper-graphs.
Next, we perform temporal pooling, and then the obtained spatial information is the evidence to determine the optimal hyper-graph construction. This can efficiently decouple the temporal and spatial clues. After performing temporal average-pooling, we can obtain $\mathbf{\bar{F}_{in}}\in\mathbb{R}^{C_{in}\times V}$.
The hyper-graph is depicted by the incidence matrix and weight matrix. So, we use a separate mapping function for each head to embed the features into the subspace for constructing the hyper-graph. This design enables preserving the original features to reflect their spatial relevance. After that, we introduce the A-NHG to obtain the incidence matrix. For the weight matrix, we adopt the Channel Attention [54]. The hyper-graph obtained operation can be formulated as follows:
$H$ $=\uplus^{8}_{i=1}\text{A-NHG}(\mathbf{\bar{F}^{k}_{in}}),$ (9)
$W$ $=\sigma^{2}(\Psi^{2}(\uplus^{8}_{i=1}\sigma^{1}(\Psi^{1}_{i}(\mathbf{\bar{F}^{k}_{in}})))),$ (10)